OpenStudy (anonymous):

find the consecutive even integers such as the first times the third is 2 greater than 5 times the second.

OpenStudy (anonymous):

hahaha.. wala pang nag re-response. xD

OpenStudy (unklerhaukus):

\[a\times(a+2)+2=5(a+1)\]

OpenStudy (anonymous):

@Diyadiya pls help ^^